.The two stars in a binary system have an orbital period of 5 years and an orbital separation, a, of 10AU. What is their combined mass?
1
Expert's answer
2020-11-23T10:31:58-0500

According to the third Kepler's law,

"\displaystyle (m+M) P^2 = a^3 \; \; \Rightarrow \; \; m+M = \frac{a^3}{P^2} = \frac{10^3}{5^2} =\frac{1000}{25} = 40" solar masses.

Answer: 40 solar masses.